A cylindrical gas tube has a pressure of 7.0 atm at 75 0C. What is the pressure if the temperature is increased to 95 0C?


1
Expert's answer
2021-10-10T10:17:56-0400

P1/T1= P2/T2

7/348= x/368

X= 7.4atm